OpenStudy (fatemahzahrah):

Show that the equation sin(x-60) - cos(30-x) = 1 can be written in the form cosx = k, where k is a constant. can somene help me with this please....

OpenStudy (atrineas):

Welcometo open study ^-^ @fatemahzahrah . . .and sorry don't know how to do this >~<

OpenStudy (egenriether):

Too long to type, wrote it out on paper

OpenStudy (mww):

expand. You have sinxcos(60) - cosxsin60 - (cos30cosx + sin30sinx) = \[\frac{ 1 }{ 2 } \sin x - \frac{ \sqrt{3} }{ 2} \cos x - \frac{ \sqrt{3} }{ 2 } \cos x - \frac{ 1 }{ 2} \sin x\] Now simplify and you should be left with cos x = k for some constant k.

OpenStudy (fatemahzahrah):

thank you @Atrineas

OpenStudy (atrineas):

No problem ^-^